# Finite-time blowup in a five- and higher-dimensional chemotaxis model with indirect signal production

### v2

@@ 4. Energy bounded below by sublinear dissipation rate
$-\mathcal{F}$ $\leq C(m,M,B,\kappa)(\rho^{2}\|\nabla f\|_{L^{2}(\Omega)}^{2}+\|\nabla f\|_{L^{2}(\Omega)}^{2n/(n+2)}+\|g\|_{L^{2}(\Omega)}+\rho^{n-2\kappa})$
holds for all $(u,v)\in\mathcal{S}$. If $\rho<R/2$, then
$-\mathcal{F}$ $\leq C(m,M,B,\kappa)(\rho^{2}\|\nabla f\|_{L^{2}(\Omega)}^{2}+2\|\nabla f\|_{L^{2}(\Omega)}^{2n/(n+2)}+\|g\|_{L^{2}(\Omega)})$ (4.27)
$\leq C(m,M,B,\kappa)\left(3\|\nabla f\|_{L^{2}(\Omega)}^{2\theta}+\|g\|_{L^{2}(\Omega)}^{2\theta}+3\right)$
$\leq C(m,M,B,\kappa)(\mathcal{D}^{\theta}+1),$
by Young inequality and concavity of the mapping $[0,\infty)\ni\eta\mapsto\eta^{\theta}\in[0,\infty)$. On the other hand, if $\rho=R/2$, then we can estimate analogously
$-\mathcal{F}$ $\leq C(m,M,B,\kappa)\left(\rho^{2}\|\nabla f\|_{L^{2}(\Omega)}^{2}+\|\nabla f\|_{L^{2}(\Omega)}^{2n/(n+2)}+\|g\|_{L^{2}(\Omega)}+\frac{2^{2\kappa-n}}{R^{2\kappa-n}}\right)$ (4.28)
$\leq C(m,M,B,\kappa)\left(2\|\nabla f\|_{L^{2}(\Omega)}^{2\theta}+\|g\|_{L^{2}(\Omega)}^{2\theta}+3+\frac{2^{2\kappa-n}}{R^{2\kappa-n}}\right)$
$\leq C(m,M,B,\kappa)(\mathcal{D}^{\theta}+1).$
Hence, (4.26) follows from (4.27) and (4.28). ∎
5. Finite-time blowup enforced by Low initial energy
This section is devoted to the proof of Theorem 1.1. For initial data $(u_{0},v_{0})$ with low energy $\mathcal{F}(u_{0},v_{0})$, we may deduce a superlinear differential inequality for the mapping $t\mapsto-\mathcal{F}(u(\cdot,t),v(\cdot,t))$ from (4.26), which implies that $(u,v)$ cannot exist globally.
@@ 5. Finite-time blowup enforced by Low initial energy
Lemma 2.3 and Proposition 3.1 give $c_{2}>0$, $c_{3}>0$ and $c_{4}=c_{4}(\kappa)>0$ such that whenever $(u_{0},v_{0})$ complies with (1.6), the corresponding solution $(u,v,w)$ satisfies $L^{1}$ estimate
$\|v_{t}\|_{L^{1}(\Omega)}\leq c_{2}(m+\|v_{0}\|_{W^{2,2}(\Omega)})\quad\text{for all }t\in(0,T_{\max}(u_{0},v_{0})),$ (5.2)
@@
$\||x|^{n-1}w\|_{W^{1,\infty}(\Omega)}\leq c_{3}m\quad\text{for all }t\in(0,T_{\max}(u_{0},v_{0})),$ (5.3)
@@
$\||x|^{\kappa-1}v\|_{W^{1,\infty}(\Omega)}\leq c_{4}(m+\|v_{0}\|_{W^{2,2}(\Omega)})$ (5.4)
@@
$B$ $:=\max\{c_{1}A,m,c_{3}m\}\quad\text{and}\quad M:=\max\{c_{2}(m+A),c_{4}(m+A)\}$
@@
$-\mathcal{F}(\tilde{u},\tilde{v})\leq c_{5}(\mathcal{D}^{\theta}(\tilde{u},\tilde{v})+1)\quad\text{for all }(\tilde{u},\tilde{v})\in\mathcal{S}(m,M,B,\kappa).$ (5.5)
@@
Indeed, we know from Proposition 2.1 that for any given $(u_{0},v_{0})\in\widetilde{\mathcal{B}}(m,A,\kappa)$, the corresponding solution $(u,v,w)$ is smooth and radially symmetric with $u>0$ and $v>0$ in $\overline{\Omega}\times[0,T_{\max}(u_{0},v_{0}))$. We infer from (5.2)–(5.4) that
@@
for all $t\in(0,T_{\max}(u_{0},v_{0}))$ by (2.4) and (2.5), it follows that $(u,v)\in\mathcal{S}(m,M,B,\kappa)$ for all $t\in(0,T_{\max}(u_{0},v_{0}))$ and thus (5.5) may be applied to $(\tilde{u},\tilde{v})=(u(\cdot,t),v(\cdot,t))$ for all $t\in(0,T_{\max}(u_{0},v_{0}))$. From (2.2) and (5.6), we see that
@@
and thus we may invert (5.5) to get
$-\frac{\;\mathrm{d}}{\;\mathrm{d}t}\mathcal{F}=\mathcal{D}\geq\left(\frac{-\mathcal{F}-c_{5}}{c_{5}}\right)^{\frac{1}{\theta}},\quad\text{for all }t\in(0,T_{\max}(u_{0},v_{0})).$ (5.7)
We shall see $T_{\max}(u_{0},v_{0})\leq T(m,A,\kappa)$. Otherwise a direct comparison argument by solving (5.7) gives
@@
We only need to fix an arbitrary $\kappa>n-2$ and apply Lemma 5.1 to find that the conclusion holds if we let $K(m,A):=K(m,A,\kappa)$ and $T(m,A):=T(m,A,\kappa)$ with $K(m,A,\kappa)$ and $T(m,A,\kappa)$ as given by Lemma 5.1. ∎
6. Initial data with large negative energy
@@
Let $\Omega=B_{R}\subset\mathbb{R}^{n}$ with some $R>0$ and $n\geq 5$. Then for any choice of a pair of positive and radially symmetric functions $(u_{0},v_{0})$ satisfying (1.6), one can find a family
$\{(u_{\eta},v_{\eta})\in L^{\infty}(\Omega)\times W^{2,\infty}(\Omega)\}_{\eta\in(0,\eta_{\star})}$
of function couples for some $\eta_{\star}\in(0,1)$ such that $(u_{\eta},v_{\eta})$ satisfies (1.6) and (1.7) for all $\eta\in(0,\eta_{\star})$, that
$\int_{\Omega}u_{\eta}=\int_{\Omega}u_{0}\quad\text{for all }\eta\in(0,\eta_{\star}),$ (6.1)
@@
$u_{\eta}\to u_{0}\quad\text{in }L^{p}(\Omega)\text{ for all }p\in\left[1,\frac{2n}{n+4}\right)\text{ as }\eta\searrow 0,$ (6.3)
@@
$\mathcal{F}(u_{\eta},v_{\eta})\to-\infty,\quad\text{as }\eta\searrow 0,$ (6.4)
where $\mathcal{F}$ is taken from (2.3).
@@
Let $\phi\in C_{0}^{\infty}(\mathbb{R}^{n})$ be a radially symmetric and nonnegative function with $\int_{\mathbb{R}^{n}}\phi=1$, which is nonincreasing with respect to $|x|$ and supported in $B_{1}$. Fixing $\gamma>1$, we define for $\eta\in(0,\eta_{\star})$,
$u_{\eta}(x):=u_{0}(x)+\left(\ln\frac{1}{\eta}\right)^{2\gamma}\eta^{-\frac{n}{2}-2}\phi\left(\frac{x}{\eta}\right)-\left(\ln\frac{1}{\eta}\right)^{2\gamma}\frac{\eta^{\frac{n}{2}-2}}{|\Omega|},\quad x\in\overline{\Omega},$
@@
$v_{\eta}(x):=v_{0}(x)+\left(\ln\frac{1}{\eta}\right)^{-\gamma}\eta^{2-\frac{n}{2}}\phi\left(\frac{x}{\eta}\right),\quad x\in\overline{\Omega},$
where $\eta_{\star}\in(0,1)\cap(0,R)$ is fixed such that
$\eta^{\frac{n}{2}-2}\left(\ln\frac{1}{\eta}\right)^{2\gamma}<|\Omega|\iota\quad\text{with }\iota:=\min_{x\in\overline{\Omega}}u_{0},\quad\text{for all }\eta\in(0,\eta_{\star}),$ (6.5)
@@
$2\eta^{\frac{n}{2}-2}\left(\ln\frac{1}{\eta}\right)^{2\gamma}<1\quad\text{for all }\eta\in(0,\eta_{\star}).$ (6.6)
A direct computation can verify the mass identity (6.1). Since
$\|v_{\eta}-v_{0}\|_{W^{2,2}(\Omega)}\leq\left(\ln\frac{1}{\eta}\right)^{-\gamma}\|\phi\|_{W^{2,2}(B_{1})}$
are valid for all $\eta\in(0,\eta_{\star})$, $\gamma>0$ ensures (6.2). We estimate by triangle inequality,
$\|u_{\eta}-u_{0}\|_{L^{p}(\Omega)}\leq\left(\ln\frac{1}{\eta}\right)^{2\gamma}\eta^{-2-\frac{n}{2}+\frac{n}{p}}\|\phi\|_{L^{p}(B_{1})}+\left(\ln\frac{1}{\eta}\right)^{2\gamma}\eta^{\frac{n}{2}-2}|\Omega|^{\frac{1}{p}-1}\quad\text{for all }\eta\in(0,\eta_{\star})$
and (6.3) follows due to $-2-n/2+n/p>0$ and $n>4$. To verify (6.4), we evaluate each term from $\mathcal{F}$. Young inequality entails
$\int_{\Omega}|v_{\eta}|^{2}\leq 2\|v_{0}\|_{L^{2}(\Omega)}^{2}+2\left(\ln\frac{1}{\eta}\right)^{-2\gamma}\eta^{4}\|\phi\|_{L^{2}(B_{1})}^{2}\quad\text{for all }\eta\in(0,\eta_{\star})$ (6.7)
and
$\int_{\Omega}|\Delta v_{\eta}|^{2}\leq 2\|\Delta v_{0}\|_{L^{2}(\Omega)}^{2}+2\left(\ln\frac{1}{\eta}\right)^{-2\gamma}\|\Delta\phi\|_{L^{2}(B_{1})}^{2}\quad\text{for all }\eta\in(0,\eta_{\star}).$ (6.8)
Thank to (6.5), $u_{\eta}$ is nonnegative for all $\eta\in(0,\eta_{\star})$, thus we can drop the nonnegative summands and estimate from below
$\int_{\Omega}u_{\eta}v_{\eta}\geq\left(\ln\frac{1}{\eta}\right)^{\gamma}\int_{\Omega}\phi^{2}\left(\frac{x}{\eta}\right)\eta^{-n}\;\mathrm{d}x=\left(\ln\frac{1}{\eta}\right)^{\gamma}\|\phi\|_{L^{2}(B_{1})}^{2}\quad\text{for all }\eta\in(0,\eta_{\star}).$ (6.9)
Dropping the possible negative part and using the monotonicity of the logarithmic function $\xi\mapsto\ln\xi$ for $\xi>0$, we compute with the help of (6.6)
$\int_{\Omega}u_{\eta}\ln u_{\eta}$ $=\int_{\Omega}u_{\eta}\ln\left(1+\frac{(-\ln\eta)^{2\gamma}\eta^{-\frac{n}{2}-2}}{u_{0}(x)}\phi\left(\frac{x}{\eta}\right)-\frac{(-\ln\eta)^{2\gamma}\eta^{\frac{n}{2}-2}}{u_{0}(x)|\Omega|}\right)+\int_{\Omega}u_{\eta}\ln u_{0}$ (6.10)
$\leq\int_{\Omega}u_{\eta}\ln\left(1+\frac{(-\ln\eta)^{2\gamma}\eta^{-\frac{n}{2}-2}}{u_{0}(x)}\phi\left(\frac{x}{\eta}\right)\right)$
$+\int_{\Omega}u_{0}\ln u_{0}+2\eta^{\frac{n}{2}-2}\left(\ln\frac{1}{\eta}\right)^{2\gamma}\|\ln u_{0}\|_{L^{\infty}(\Omega)}$
$\leq\int_{\Omega}u_{0}\ln u_{0}+\|\ln u_{0}\|_{L^{\infty}(\Omega)}+m\ln\left(1+\left(\ln\frac{1}{\eta}\right)^{2\gamma}\frac{\phi(0)}{\eta^{\frac{n}{2}+2}\iota}\right)$
for all $\eta\in(0,\eta_{\star})$. Using (6.6), we calculate
$\ln\left(1+\left(\ln\frac{1}{\eta}\right)^{2\gamma}\frac{\phi(0)}{\eta^{\frac{n}{2}+2}\iota}\right)$ (6.11)
$=2\gamma\ln\ln\frac{1}{\eta}+\frac{n+4}{2}\ln\frac{1}{\eta}+\ln\frac{\phi(0)}{\iota}+\ln\left(1+\left(\ln\frac{1}{\eta}\right)^{-2\gamma}\frac{\eta^{\frac{n}{2}+2}\iota}{\phi(0)}\right)$
$\leq n\ln\frac{1}{\eta}+\ln\frac{\phi(0)}{\iota}+\ln\left(1+\frac{\iota}{\phi(0)}\right)=n\ln\frac{1}{\eta}+\ln\left(1+\frac{\phi(0)}{\iota}\right)$
for all $\eta\in(0,\eta_{\star})$. Collecting (6.7)–(6.11) gives
$\mathcal{F}(u_{\eta},v_{\eta})=\int_{\Omega}(u_{\eta}\log u_{\eta}-u_{\eta}v_{\eta})+\frac{1}{2}\int_{\Omega}|(-\Delta+1)v_{\eta}|^{2}$ (6.12)
$=-\left(\ln\frac{1}{\eta}\right)^{\gamma}\|\phi\|_{L^{2}(B_{1})}^{2}+2\left(\ln\frac{1}{\eta}\right)^{-2\gamma}(\|\Delta\phi\|_{L^{2}(B_{1})}^{2}+\|\phi\|_{L^{2}(B_{1})}^{2})+nm\ln\frac{1}{\eta}$
$+2\|\Delta v_{0}\|_{L^{2}(\Omega)}^{2}+2\|v_{0}\|_{L^{2}(\Omega)}^{2}+\int_{\Omega}u_{0}\ln u_{0}+\|\ln u_{0}\|_{L^{\infty}(\Omega)}+m\ln\left(1+\frac{\phi(0)}{\iota}\right),$
for all $\eta\in(0,\eta_{\star})$. In view of $\gamma>1$, we have $\mathcal{F}(u_{\eta},v_{\eta})\to-\infty$ as $\eta\searrow 0$ and complete the proof. ∎
@@
Let initial datum $(u_{0},v_{0})$ be specified as in Theorem 1.2. Then the family $\{(u_{\eta},v_{\eta})\}_{\eta\in(0,\eta_{\star})}$ constructed in Lemma 6.1 clearly satisfies two limit formulae according to (6.2)–(6.4).
Let us now fix $\varepsilon>0$, $\|u_{0}\|_{L^{1}(\Omega)}=m$ and $A=\|v_{0}\|_{W^{2,2}(\Omega)}$ given as in Theorem 1.2. We may infer the existence of positive constants $K(m,A+\varepsilon)$ and $T(m,A+\varepsilon)$ from Theorem 1.1 such that for any $(\tilde{u}_{0},\tilde{v}_{0})$ from the set $\mathcal{B}(m,A+\varepsilon)$ defined by (1.8), the corresponding classical solution with maximal existence time $T_{\max}(\tilde{u}_{0},\tilde{v}_{0})$ given by Proposition 2.1, blows up in finite time $T_{\max}(\tilde{u}_{0},\tilde{v}_{0})\leq T(m,A+\varepsilon)$. Thank to (6.1), (6.2) and (6.4), we may pick some $\eta_{\star\star}\in(0,\eta_{\star})$ such that $(u_{\eta},v_{\eta})\in\mathcal{B}(m,A+\varepsilon)$ is valid for all $\eta\in(0,\eta_{\star\star})$. ∎
Acknowledgments
Supported in part by National Natural Science Foundation of China (No. 12271092, No. 11671079) and the Jiangsu Provincial Scientific Research Center of Applied Mathematics (No. BK20233002).

### v3

@@ 4 Energy bounded below by sublinear dissipation rate
-$\mathcal{F}$ $\leq C(m,M,B,\kappa)(\rho^{2}\|\nabla f\|_{L^{2}(\Omega)}^{2}+\|\nabla f\|_{L^{2}(\Omega)}^{2n/(n+2)}+\|g\|_{L^{2}(\Omega)}+\rho^{n-2\kappa})$
holds for all $(u,v)\in\mathcal{S}(m,M,B,\kappa)$. If $\rho<R/2$, then
$\begin{split}-\mathcal{F}&\leq C(m,M,B,\kappa)(\rho^{2}\|\nabla f\|_{L^{2}(\Omega)}^{2}+\|\nabla f\|_{L^{2}(\Omega)}^{2n/(n+2)}+\|g\|_{L^{2}(\Omega)}+\|\nabla f\|_{L^{2}(\Omega)}^{n/(n+2)})\\ &\leq C(m,M,B,\kappa)\left(3\|\nabla f\|_{L^{2}(\Omega)}^{2\theta}+\|g\|_{L^{2}(\Omega)}^{2\theta}+4\right)\\ &\leq C(m,M,B,\kappa)(\mathcal{D}^{\theta}+1),\end{split}$ (48)
by Young inequality and monotonicity of the mapping $0,\infty)\ni\eta\mapsto\eta^{\theta}\in[0,\infty)$. On the other hand, if $\rho=R/2$, then we can estimate analogously
$\begin{split}-\mathcal{F}&\leq C(m,M,B,\kappa)\left(\rho^{2}\|\nabla f\|_{L^{2}(\Omega)}^{2}+\|\nabla f\|_{L^{2}(\Omega)}^{2n/(n+2)}+\|g\|_{L^{2}(\Omega)}+\frac{2^{2\kappa-n}}{R^{2\kappa-n}}\right)\\ &\leq C(m,M,B,\kappa)\left(2\|\nabla f\|_{L^{2}(\Omega)}^{2\theta}+\|g\|_{L^{2}(\Omega)}^{2\theta}+3+\frac{2^{2\kappa-n}}{R^{2\kappa-n}}\right)\\ &\leq C(m,M,B,\kappa)(\mathcal{D}^{\theta}+1).\end{split}$ (49)
Hence, (47) follows from (48) and (49). ∎
5 Finite-time blowup enforced by low initial energy
This section is devoted to the proof of Theorem 1.1. For initial data $(u_{0},v_{0})$ with low energy $\mathcal{F}(u_{0},v_{0})$, we may deduce a superlinear differential inequality for the mapping $t\mapsto-\mathcal{F}(u(\cdot,t),v(\cdot,t))$ from (47), which implies that $(u,v)$ cannot exist globally.
@@ 5 Finite-time blowup enforced by low initial energy
Lemma 2.3 and Proposition 3.1 give $c_{2}>0$, $c_{3}>0$ and $c_{4}=c_{4}(\kappa)>0$ such that whenever $(u_{0},v_{0})$ complies with (6), the corresponding solution $(u,v,w)$ satisfies $L^{1}$ estimate
$\|v_{t}\|_{L^{1}(\Omega)}\leq c_{2}(m+\|v_{0}\|_{W^{2,2}(\Omega)})\quad\text{for all }t\in(0,T_{\max}(u_{0},v_{0})),$ (51)
@@
$\||x|^{n-1}w\|_{W^{1,\infty}(\Omega)}\leq c_{3}m\quad\text{for all }t\in(0,T_{\max}(u_{0},v_{0})),$ (52)
@@
$\||x|^{\kappa-1}v\|_{W^{1,\infty}(\Omega)}\leq c_{4}(m+\|v_{0}\|_{W^{2,2}(\Omega)})$ (53)
@@
$B$ $:=\max\{c_{1}A,m,c_{3}m\}\quad\text{and}\quad M:=\max\{c_{2}(m+A),c_{4}(m+A)\},$
@@
[…]
@@
Indeed, we know from Proposition 2.1 that for any given $(u_{0},v_{0})\in\widetilde{\mathcal{B}}(m,A,\kappa)$, the corresponding solution $(u,v,w)$ is smooth and radially symmetric with $u>0$ and $v>0$ in $\overline{\Omega}\times[0,T_{\max}(u_{0},v_{0}))$. We infer from (51)–(53) that
@@
for all $t\in(0,T_{\max}(u_{0},v_{0}))$ by (12) and (13), it follows that $(u,v)\in\mathcal{S}(m,M,B,\kappa)$ for all $t\in(0,T_{\max}(u_{0},v_{0}))$ and thus (54) may be applied to $(\tilde{u},\tilde{v})=(u(\cdot,t),v(\cdot,t))$ for all $t\in(0,T_{\max}(u_{0},v_{0}))$. From (10) and (55), we see that
@@
and thus we may invert (54) to get
[…]
We shall see $T_{\max}(u_{0},v_{0})\leq T(m,A,\kappa)$. Otherwise a direct comparison argument by solving (56) gives
@@
We only need to fix an arbitrary $\kappa>n-2$ and apply Lemma 5.1 to find that the conclusion holds if we let $K(m,A):=K(m,A,\kappa)$ and $T(m,A):=T(m,A,\kappa)$ with $K(m,A,\kappa)$ and $T(m,A,\kappa)$ given by Lemma 5.1. ∎
6 Initial data with large negative energy
@@
Let $n\geq 5$ and $\Omega\subset\mathbb{R}^{n}$ be a bounded and smooth domain. Then for any choice of function pair $(u_{0},v_{0})$ satisfying (6), one can find a sequence
$\{(u_{k},v_{k})\in L^{\infty}(\Omega)\times W^{2,\infty}(\Omega)\}_{k\in\mathbb{N}}$
of function couples such that $(u_{k},v_{k})$ satisfies (6) for all $k\in\mathbb{N}$, that
$\int_{\Omega}u_{k}=\int_{\Omega}u_{0}\quad\text{for all }k\in\mathbb{N},$ (57)
@@
[…]
@@
$\mathcal{F}(u_{k},v_{k})\to-\infty,\quad\text{as }k\to 0,$ (60)
where $\mathcal{F}$ is taken from (11).
@@
Without loss of generality, we may assume that $B_{R}\subset\Omega$ for some $R>0$. Let $\phi\in C_{0}^{\infty}(\mathbb{R}^{n})$ be a radially symmetric and nonnegative function with $\int_{\mathbb{R}^{n}}\phi=1$, which is compactly supported in $B_{1}$. Fixing $\gamma>0$, we define for $k\in\mathbb{N}$,
$u_{k}(x):=u_{0}(x)+\left(\ln\frac{1}{\eta}\right)^{2\gamma}\eta^{-\frac{n}{2}-2}\phi\left(\frac{x}{\eta}\right)-\left(\ln\frac{1}{\eta}\right)^{2\gamma}\eta^{\frac{n}{2}-2}\cdot\frac{u_{0}(x)}{\int_{\Omega}u_{0}},\quad x\in\overline{\Omega},$
@@
$v_{k}(x):=v_{0}(x)+\left(\ln\frac{1}{\eta}\right)^{-\gamma}\eta^{2-\frac{n}{2}}\phi\left(\frac{x}{\eta}\right),\quad x\in\overline{\Omega},$
where
$\eta=\eta(k):=\frac{\eta_{\star}}{k+1},\quad k\in\mathbb{N},$
and $\eta_{\star}\in(0,1)\cap(0,R)$ is fixed such that
$\eta^{\frac{n}{2}-2}\left(\ln\frac{1}{\eta}\right)^{2\gamma}<\int_{\Omega}u_{0}\quad\text{for all }k\in\mathbb{N}.$ (61)
A direct computation can verify the mass identity (57). Since
$\|v_{k}-v_{0}\|_{W^{2,2}(\Omega)}\leq\left(\ln\frac{1}{\eta}\right)^{-\gamma}\|\phi\|_{W^{2,2}(B_{1})}$
are valid for all $k\in\mathbb{N}$, $\gamma>0$ ensures (58). We estimate by triangle inequality,
$\begin{split}\|u_{k}-u_{0}\|_{L^{p}(\Omega)}&\leq\left(\ln\frac{1}{\eta}\right)^{2\gamma}\eta^{-2-\frac{n}{2}+\frac{n}{p}}\|\phi\|_{L^{p}(B_{1})}\\ &\quad+\left(\ln\frac{1}{\eta}\right)^{2\gamma}\eta^{\frac{n}{2}-2}\cdot\frac{\|u_{0}\|_{L^{p}(\Omega)}}{\int_{\Omega}u_{0}}\quad\text{for all }k\in\mathbb{N}\text{ and }p\in[1,\infty).\end{split}$
Then (59) follows due to 2n/2+n/p>0-2-n/2+n/p>0 and $n>4$. To verify (60), we evaluate each term from $\mathcal{F}$. Young inequality entails
$\int_{\Omega}|v_{k}|^{2}\leq 2\|v_{0}\|_{L^{2}(\Omega)}^{2}+2\left(\ln\frac{1}{\eta}\right)^{-2\gamma}\eta^{4}\|\phi\|_{L^{2}(B_{1})}^{2}\quad\text{for all }k\in\mathbb{N}$ (62)
@@
$\int_{\Omega}|\Delta v_{k}|^{2}\leq 2\|\Delta v_{0}\|_{L^{2}(\Omega)}^{2}+2\left(\ln\frac{1}{\eta}\right)^{-2\gamma}\|\Delta\phi\|_{L^{2}(B_{1})}^{2}\quad\text{for all }k\in\mathbb{N}.$ (63)
Thanks to (61), $u_{k}$ is nonnegative for all $k\in\mathbb{N}$, thus we can drop the nonnegative summands and estimate from below
$\begin{split}\int_{\Omega}u_{k}v_{k}&\geq\left(\ln\frac{1}{\eta}\right)^{\gamma}\int_{\Omega}\phi^{2}\left(\frac{x}{\eta}\right)\eta^{-n}\;\mathrm{d}x\\ &=\left(\ln\frac{1}{\eta}\right)^{\gamma}\|\phi\|_{L^{2}(B_{1})}^{2}\quad\text{for all }k\in\mathbb{N}.\end{split}$ (64)
Noting that
$\xi^{q}-\xi\geq(q-1)\xi\ln\xi\quad\text{for all }\xi\in(0,\infty)\text{ and }q\in(1,\infty),$
we compute
$\begin{split}\int_{\Omega}u_{k}\ln u_{k}&\leq\int_{\Omega}\frac{u_{k}^{q}-u_{k}}{q-1}\leq\int_{\Omega}\frac{u_{k}^{q}}{q-1}\\ &\leq\frac{2^{q-1}}{q-1}\bigg(\|u_{0}\|_{L^{q}(\Omega)}^{q}+\Big(\ln\frac{1}{\eta}\Big)^{2q\gamma}\eta^{-\frac{nq}{2}-2q+n}\|\phi\|_{L^{q}(B_{1})}^{q}\bigg)\end{split}$ (65)
for all $k\in\mathbb{N}$ and $q\in(1,\infty)$. Collecting (62)–(65) gives
$\begin{split}\mathcal{F}(u_{k},v_{k})&=\int_{\Omega}(u_{k}\log u_{k}-u_{k}v_{k})+\frac{1}{2}\int_{\Omega}|(-\Delta+1)v_{k}|^{2}\\ &\leq-\left(\ln\frac{1}{\eta}\right)^{\gamma}\|\phi\|_{L^{2}(B_{1})}^{2}+2\left(\ln\frac{1}{\eta}\right)^{-2\gamma}(\|\Delta\phi\|_{L^{2}(B_{1})}^{2}+\|\phi\|_{L^{2}(B_{1})}^{2})\\ &\quad+\frac{2^{q-1}}{q-1}\bigg(\ln\frac{1}{\eta}\bigg)^{2q\gamma}\eta^{-\frac{nq}{2}-2q+n}\|\phi\|_{L^{q}(B_{1})}^{q}\\ &\quad+2\|\Delta v_{0}\|_{L^{2}(\Omega)}^{2}+2\|v_{0}\|_{L^{2}(\Omega)}^{2}+\frac{2^{q-1}}{q-1}\|u_{0}\|_{L^{q}(\Omega)}^{q}\end{split}$
for all $k\in\mathbb{N}$ and
$q\in\bigg(1,\frac{2n}{n+4}\bigg).$
In view of $\gamma>0$, we have $\mathcal{F}(u_{k},v_{k})\to$-$\infty$ as $k\to\infty$ and complete the proof. ∎
@@
Let initial datum $(u_{0},v_{0})$ be specified as in Theorem 1.2. Then the sequence $\{(u_{k},v_{k})\}_{k\in\mathbb{N}}$ of positive function pairs constructed in Lemma 6.1 clearly satisfies the desired properties according to (57)–(60). ∎
[…]
*Acknowledgements The authors thank the anonymous referees for their helpful comments and suggestions, which greatly improve the presentation of our paper.
[…]
*Funding The first author has been supported by “the Fundamental Research Funds for the Central Universities” (No. B250201215). The second author has been supported in part by National Natural Science Foundation of China (No. 12271092, No. 11671079) and the Jiangsu Provincial Scientific Research Center of Applied Mathematics under Grant No. BK20233002.
[…]
*Conflict of interest
The authors have not disclosed any conflict of interest.